# Evaluating multiple models using labeled and unlabeled data

### v2

@@ Abstract
It remains difficult to evaluate machine learning classifiers in the absence of a large, labeled dataset. While labeled data can be prohibitively expensive or impossible to obtain, unlabeled data is plentiful. Here, we introduce Semi-Supervised Model Evaluation (SSME), a method that uses both labeled and unlabeled data to evaluate machine learning classifiers. SSME is the first evaluation method to take advantage of the fact that: (i) there are frequently multiple classifiers for the same task, (ii) continuous classifier scores are often available for all classes, and (iii) unlabeled data is often far more plentiful than labeled data. The key idea is to use a semi-supervised mixture model to estimate the joint distribution of ground truth labels and classifier predictions. We can then use this model to estimate any metric that is a function of classifier scores and ground truth labels (e.g., accuracy or expected calibration error). We present experiments in four domains where obtaining large labeled datasets is often impractical: (1) healthcare, (2) content moderation, (3) molecular property prediction, and (4) image annotation. Our results demonstrate that SSME estimates performance more accurately than do competing methods, reducing error by 5.1$\times$ relative to using labeled data alone and 2.4$\times$ relative to the next best competing method. SSME also improves accuracy when evaluating performance across subsets of the test distribution (e.g., specific demographic subgroups) and when evaluating the performance of language models.
11footnotetext: *Denotes equal contribution22footnotetext: **Denotes joint last authorship
@@ 1 Introduction
Rigorous evaluation is essential to the safe deployment of machine learning classifiers. The standard approach is to measure classifier performance using a large labeled dataset. In practice, however, labeled data is often scarce (Culotta & McCallum, 2005; Dutta & Das, 2023). Exacerbating the challenge of evaluation, the number of off-the-shelf classifiers has increased dramatically through the widespread usage of model hubs. The modern machine learning practitioner thus has a myriad of trained models, but little labeled data with which to evaluate them.
In many domains, unlabeled data is much more abundant than labeled data (Bepler et al., 2019; Sagawa et al., 2021; Movva et al., 2024). To take advantage of this, we introduce Semi-Supervised Model Evaluation (SSME), a method that can be used to evaluate multiple classifiers using both labeled and unlabeled data. Our key idea is to estimate the joint distribution of ground truth classes $y$ and continuous classifier scores $\mathbf{s}$ (i.e., transformed probabilistic predictions) using a mixture model, where different components of the mixture model correspond to different classes. The joint distribution allows us to evaluate performance on examples where we have access only to each classifier’s scores, i.e. unlabeled examples. SSME can estimate any metric which is a function of class labels and probabilistic predictions, which includes widely-used metrics like accuracy, expected calibration error, AUC, and AUPRC.
SSME is the first evaluation method to capture three key facets of modern machine learning settings: (i) multiple machine learning classifiers, (ii) probabilistic predictions over all classes, and (iii) unlabeled data.
Simultaneously using all three is difficult because it requires accurately estimating the (potentially high-dimensional) joint distribution $P(y,\mathbf{s})$ with primarily unlabeled data.
While prior work captures subsets of these properties (Welinder et al., 2013; Platanios et al., 2017; Ji et al., 2020; Chouldechova et al., 2022; Boyeau et al., 2024) — for example, augmenting labeled data with unlabeled data to evaluate a single classifier — no existing approach accommodates all three.
We show that using all available data — multiple classifiers, continuous scores over all classes, and unlabeled data — enables SSME to produce more accurate performance estimates compared to prior work. We test SSME’s ability to estimate the absolute performance of each classifier across eight tasks, four metrics, and dozens of classifiers, where SSME accepts a set of classifiers, little labeled data (i.e. between 20 and 100 labeled examples) and more abundant unlabeled data (i.e. 1000 unlabeled examples).
Concretely, we make four contributions:
@@
    We propose SSME, a method to evaluate multiple classifiers using labeled and unlabeled data. SSME can be applied to any number of classifiers and classes and is able to estimate any metric that compares predicted probabilities to ground truth labels.
@@
    We conduct semi-synthetic experiments to characterize factors affecting SSME’s performance: the accuracy, calibration, and cardinality of the classifier set being evaluated.
@@
    We show, in experiments spanning multiple data modalities, domains, and classifier architectures, that SSME achieves the lowest metric estimation error compared to using labeled data alone and compared to prior work, across all considered metrics.
  4. 4.
    We validate two broadly useful applications of SSME: evaluating subgroup-specific performance, a critical step in assessing algorithmic fairness, and evaluating language models.
2 Related Work
Our work builds on two areas of literature: methods that use a combination of labeled and unlabeled data to 1) evaluate a single classifier, or 2) evaluate the accuracy of multiple discrete annotations. For a discussion of connections to prediction-powered inference and unsupervised classifier evaluation, see Appendix A.
Semi-supervised evaluation of single classifiers involves the evaluation of a single classifier using both labeled and unlabeled data. There are two types of assumptions common in this literature. The first places parametric constraints on the distribution of classifier scores. Several works attempt to fit a mixture model to the distribution of classifier scores (Welinder et al., 2013; Chouldechova et al., 2022; Miller et al., 2018), as we do, while others apply techniques from Bayesian calibration (Ji et al., 2020; 2021). Our work differs in that the proposed framework naturally capitalizes on multiple classifiers, and as our results show, doing so results in improved estimates of performance. The second type of assumption relates to the structure of the shift between the labeled and unlabeled data; as Garg et al. (2022) establish, estimating accuracy on the unlabeled data is impossible absent assumptions about the nature of the distribution shift.
Semi-supervised evaluation of discrete annotators was first introduced by Dawid & Skene (1979), who proposed a method to estimate ground truth in the presence of multiple potentially noisy discrete annotations. Many follow-on works inherit Dawid-Skene’s strong assumption of class-conditional independence of annotator errors (Parisi et al., 2014; Platanios et al., 2017), including popular approaches in weak supervision (Ratner et al., 2017; Bach et al., 2017; Fu et al., 2020), where annotators are instead user-provided labeling functions. Such an assumption is plausible in certain contexts, but does not naturally translate to sets of candidate classifiers, whose predictions are likely to be correlated. Subsequent work has made an effort to relax the assumption of class-conditional independence, replacing it with independence conditional on a latent notion of example difficulty (Paun et al., 2018) or adjusting for dependencies between annotators (Ratner et al., 2017; Bach et al., 2017; Fu et al., 2020). However, these methods are designed to estimate the accuracy of binary annotations; they do not exploit the continuous probabilities available in multi-classifier evaluation. While some work has made progress towards accommodating continuous predicted probabilities (Nazabal et al., 2016; Pirš & Štrumbelj, 2019), their focus is optimal aggregation, in contrast to our own, which is evaluation. Recent work uses a continuous notion of classifier confidence in conjunction with discrete annotations from each annotator (Goh et al., 2022; Boyeau et al., 2024), but do not use the distribution of classifier scores over all classes.
Figure 1: Using SSME with two binary classifiers. (1) Obtain classifier scores on all examples, a small subset of which are labeled. (2) Fit a mixture model to estimate the joint density of scores and labels (where $s_{1}$ corresponds to the score assigned by the first classifier, and $s_{2}$ the second). (3) Use the resulting density to estimate metrics such as accuracy or expected calibration error. For instance, for the unlabeled blue point in panel 2, we sample a label $y$ several times (left); if the example were labeled (right), we would use the true $y$. SSME can be applied to any number of classifiers and classes $K$, and supports any metric that compares classifier scores to ground truth labels.
@@ 3 Problem Setting
We consider a setting in which a practitioner wishes to evaluate several classifiers. Formally, there are $M$ classifiers $[f_{1},f_{2},\ldots,f_{M}]$ designed for the same task. These classifiers may differ in their training data, function class, or training hyperparameters, among other possibilities.
Each classifier in the set maps from the same input domain $\mathcal{X}$ to a probability distribution over $K$ classes, i.e. $f_{j}:\mathcal{X}\rightarrow\Delta^{K-1}$. Let $\mathbf{s}^{(i)}=[f_{1}(x^{(i)}),f_{2}(x^{(i)}),\ldots,f_{M}(x^{(i)})]$ denote the concatenated set of classifier scores on a particular data point $x^{(i)}$.
During evaluation, we have access to the set of classifiers and two datasets: (1) a small labeled dataset, $\mathcal{D}_{L}=\{(x^{(i)},y^{(i)})\}_{i=1}^{n_{\ell}}$ and (2) a larger unlabeled dataset $\mathcal{D}_{U}=\{(x^{(i)})\}_{i=1}^{n_{u}}$. The goal is to estimate classifier performance using standard metrics such as expected calibration error (ECE) or accuracy. If one knew the true label $y^{(i)}$ for each point $x^{(i)}$, it would be straightforward to evaluate the performance of each pre-trained classifier. However, in practice, the true label is not available for unlabeled examples, so we aim to infer a distribution over ground truth labels. We assume that unlabeled data is far more available than labeled data, i.e. $n_{u}>>n_{\ell}$.
@@ 4 Method
Our aim is to develop an approach that captures three common properties of modern classification settings: (i) multiple available trained models, (ii) an abundance of unlabeled data, relative to labeled data, and (iii) access to each classifier’s predicted class probabilities on every input.
The core idea underlying SSME is that if we can estimate the joint distribution of ground truth labels $y$ and classifier scores $\mathbf{s}$ — i.e., $P(y,\mathbf{s})$ — then we can estimate any metric that is a function of classifier scores and ground truth labels. Notably, we use the classifier scores $\mathbf{s}$ (i.e. the probability outputs over all classes) and not the raw inputs $\mathbf{x}$ to estimate this joint density, consistent with prior work (Ji et al., 2020; Boyeau et al., 2024). The classifier scores $\mathbf{s}$ and the corresponding labels $y$ are sufficient to characterize most standard metrics, including calibration and accuracy. Given an estimate for the joint distribution over $(y,\mathbf{s})$, one could estimate the accuracy of the first model $f_{1}$ by (repeatedly) drawing labels $\hat{y}$ from $P(y|\mathbf{s})$ for each data point $\mathbf{s}$ and compute agreement between the drawn label and the first model’s predicted label. We estimate this joint density $P(y,\mathbf{s})$ using a mixture model, which enables us to capture class-specific variation in model scores — for instance, classifiers in our set may agree in their predictions for one class but not another.
We assume that the unlabeled samples $\mathbf{s}^{(i)}$ are drawn from the same distribution as the labeled samples $(y^{(i)},\mathbf{s}^{(i)})$. Furthermore, given the challenges of density estimation in high-dimensional settings (Wang & Scott, 2019; Rippel & Adams, 2013), we primarily focus on settings with a limited ($\leq$ 50) number of classes. Our latent variable model can support higher-dimensional evaluation problems but would require robust semi-supervised density estimation procedures.
To estimate the joint distribution $P(y,\mathbf{s})$, we maximize the log likelihood over both the labeled and unlabeled datasets. When $y$ is unobserved, we treat it as a latent variable and marginalize it out. Our latent variable model and estimation procedure is detailed in Fig. 1. Overall, we maximize the following expression, where $n_{\ell}$ and $n_{u}$ refer to the number of labeled examples and unlabeled examples, respectively:
@@
where $\lambda_{L}$ modulates the relative weight of the labeled data in the likelihood. We fix $\lambda_{L}=1$ in our main experiments; our framework can also accommodate alternative weights if hyperparameter tuning or domain knowledge justify them.
Parametrization
Our approach, SSME, can accommodate multiple parameterizations of the class-conditional distribution of scores, as long as they can be learned in the semi-supervised setting described above. We denote the parameterized distribution as $P_{\theta}(\mathbf{s}|y)$. Here, we use a kernel density estimator (KDE) to parameterize $P_{\theta}(\mathbf{s}|y)$. The KDE for the $k$th class can be written as:
$P_{\theta}(\mathbf{s}|y=k)=\frac{1}{\sum_{i}P(y^{(i)}=k|s^{(i)})}\sum_{i}\mathcal{K}_{h}(\mathbf{s}-s^{(i)})\cdot P(y^{(i)}=k|s^{(i)})$
which are parametrized by bandwidth $h$ and the kernel type $\mathcal{K}$. We weight each point by the probability it belongs to the cluster (i.e., a soft label) since we fit a separate KDE for each class. When $\mathbf{s}^{(i)}$ is labeled, $P(y^{(i)}=k)=1$ for the true label $k$, and when $\mathbf{s}^{(i)}$ is unlabeled, we assign soft labels to each point. Kernel density estimators are well-suited for the task, as they do not make parametric assumptions on the distributional form of each component. This is useful for modeling distributions of predictions, which can vary widely across outcomes and models.
Fitting densities on simplices (or more broadly bounded domains) can results in biased estimates near the boundaries, a problem known as the boundary bias problem (Jones, 1993). To overcome this challenge, we utilize invertible compositional data transforms, the preferred method for analyzing data with unit-sum constraints (Aitchison, 1982; Pawlowsky-Glahn & Buccianti, 2011). To implement this, we transform probabilistic predictions over $K$ classes to points (“scores”) in $\mathbb{R}^{K-1}$ using the additive log ratio transform, which produces a one-to-one mapping between the two spaces. Additive log-ratio transforms are a reparameterization trick; they transform the classifier scores from a hard-to-model space — the probability simplex — to an easier-to-model space — unbounded reals. For details on the additive log-ratio transform, see Appendix E.1.
Model estimation
For all experiments, we fit each kernel density estimator with a Gaussian kernel and estimate $h$ using the improved Sheather-Jones algorithm (Botev et al., 2010).
In accordance with previous semi-supervised mixture models, we fit our mixture model using expectation-maximization (EM) (Dempster et al., 1977; Zhao et al., 2023), which alternates between (1) the E-step, which estimates which mixture component each data point belongs to, and (2) the M-step, which estimates the parameters of each component based on the soft component assignments. We take the same approach, where we alternate between estimating the true label for a given example, and estimating the class-conditional distribution of model predictions based on these estimated labels. We optimize the parameters using EM over 50 epochs. We initialize component assignments by drawing a label for a given example according to the mean classifier score across the set of classifiers.
Evaluation Once we have estimated the parameters $\theta$ for our mixture model, we can use our fitted density to estimate metrics of interest, using the procedure described in the right panel of Figure 1. In particular, because we estimate the conditional distribution $P_{\theta}(y|\mathbf{s})$, for each unlabeled example (where we observe only $\mathbf{s}^{(i)}$) we can repeatedly draw ground truth labels $y$ from the distribution $P_{\theta}(y|\mathbf{s})$. To estimate the accuracy of classifier $j$, we would then compute, in expectation, how often $\mathbf{s}^{(i)}_{j}$ agrees with inferred $y^{(i)}$, and average this over all examples $i$. For unlabeled examples, we use the inferred distribution over $y^{(i)}$, and for labeled examples we use the true $y^{(i)}$. Details on our metric evaluation procedure are available in Appendix E.2.
Alternative parameterizations
SSME can also accommodate alternate parameterizations of $P(\mathbf{s}^{(i)}|y^{(i)})$ provided they can (1) accommodate both labeled and unlabeled data and (2) be fit using the mixture model framework described above. One alternative that we explore is using a normalizing flow to model each mixture component; we detail our approach and investigate when this improves over the KDE parameterization in Appendix C. While normalizing flows can also be learned in a semi-supervised setting, they often struggle to model multimodal distributions (Stimper et al., 2022; Cornish et al., 2020). We use the KDE to generate our results in binary datasets. We additionally test the normalizing flow in multiclass settings and show it performs competitively.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Large, labeled datasets are critical for evaluation in machine learning, but such datasets are often prohibitively expensive or impossible to obtain (culotta2005reducing, ; dutta2023remote, ). Exacerbating the challenge of evaluation, the number of off-the-shelf classifiers has increased dramatically through the widespread usage of model hubs. The modern machine learning practitioner thus has a myriad of trained models, but little labeled data with which to evaluate them.
In many domains, unlabeled data is much more abundant than labeled data (bepler2019positive, ; sagawa2021extending, ; movva2024using, ). To take advantage of this, we introduce Semi-Supervised Model Evaluation (SSME), a method that can be used to evaluate multiple classifiers using both labeled and unlabeled data. The key idea is to estimate the joint distribution of ground truth classes $y$ and continuous classifier scores $\mathbf{s}$ using a mixture model, where different components of the mixture model correspond to different classes. The joint distribution allows us to evaluate performance on examples where we have access only to each classifier’s scores, i.e. unlabeled examples. SSME can estimate any metric that is a function of class labels and probabilistic predictions, including widely-used metrics like accuracy, expected calibration error, AUC, and AUPRC. Concretely, we can evaluate a classifier on an unlabeled point by using our estimate of $P(y,\mathbf{s})$.
SSME is the first evaluation method to capture three key facets of modern machine learning settings: (i) multiple machine learning classifiers, (ii) probabilistic predictions over all classes, and (iii) unlabeled data. Simultaneously using all three is difficult because it requires accurately estimating the (potentially high-dimensional) joint distribution $P(y,\mathbf{s})$ with primarily unlabeled data. While prior work captures subsets of these three sources of information (welinder_lazy_2013, ; platanios_estimating_2017, ; ji_can_2020, ; chouldechova_unsupervised_2022, ; boyeau2024autoeval, ) — for example, augmenting labeled data with unlabeled data to evaluate a single classifier — no existing approach accommodates all three.
We show both theoretically and empirically that by using all available data — multiple classifiers, continuous scores over all classes, and unlabeled data — SSME is able to produce more accurate performance estimates compared to prior work. We make three contributions:
@@
    We propose SSME, a method that incorporates unlabeled data, continuous classifier scores, and multiple classifiers to produce more accurate performance estimates than prior work. SSME is able to estimate any metric that compares predicted probabilities to ground truth labels and accommodates varying numbers of classifiers and classes.
@@
    We prove that using unlabeled data and multiple classifiers improves performance estimation, providing insight into SSME’s strong empirical performance across domains.
@@
[…]
Figure 1: Using SSME with two binary classifiers. SSME applies to settings in which classifier scores are available for a set of examples, some labeled (left). SSME first fits a mixture model to estimate the joint density of scores and labels (where $s_{1}$ is the score assigned by the first classifier, and $s_{2}$ the second, middle panel). SSME next uses the mixture model to compute metrics with respect to the expected label for each example (right). SSME can be applied to any number of classifiers and classes $K$, and supports any metric that compares classifier scores to ground truth labels.
@@ 2 Problem Setting
We consider a setting in which a practitioner wishes to evaluate several classifiers using both labeled and unlabeled data. Formally, there are $M$ classifiers $f_{1},f_{2},\ldots,f_{M}]$ designed for the same task, which might differ in their training data, architectures, or optimization, among other possibilities. Each classifier maps from the same input domain $\mathcal{X}$ to a probability distribution over $K$ classes. Let $\mathbf{s}^{i}=[f_{1}(x^{i}),f_{2}(x^{i}),\ldots,f_{M}(x^{i})]$ denote the concatenated set of classifier scores for data point $x^{i}$; Figure 1 depicts the setting.
We assume access to the set of classifiers and two datasets: (1) a small labeled dataset, $\mathcal{D}_{L}=\{(x^{i},y^{i})\}_{i=1}^{n_{\ell}}$ and (2) a larger unlabeled dataset $\mathcal{D}_{U}=\{(x^{i})\}_{i=1}^{n_{u}}$. We assume that unlabeled data is far more available than labeled data, i.e. $n_{u}>>n_{\ell}$, and that the unlabeled samples are drawn from the same distribution as the labeled samples. Our goal is to use these two datasets to estimate classifier performance using standard metrics such as expected calibration error (ECE) or accuracy. If we knew the true label $y^{i}$ for each point $x^{i}$, it would be straightforward to evaluate the performance of each pre-trained classifier. However, the true label is not available for unlabeled examples, so we instead aim to infer a distribution over ground truth labels.
@@ 3 Method
Semi-supervised model evaluation (SSME) contains two key steps. First, SSME estimates the joint distribution of ground truth labels $y$ and classifier scores $\mathbf{s}$ — i.e. $P(y,\mathbf{s})$ – using both labeled and unlabeled data. Second, SSME uses the estimate of $P(y,\mathbf{s})$ to compute metrics with respect to the posterior over $y$ implied by classifier scores $\mathbf{s}$, allowing classifier evaluation using both labeled data (where ground truth is available) and unlabeled data (where ground truth is inferred).
Step 1: Estimating $P(y,\mathbf{s})$.
SSME estimates $P(y,\mathbf{s})$ using a semi-supervised mixture model, where components of the mixture model correspond to different ground truth classes. Following standard procedures in mixture model estimation, we optimize parameters of the mixture model $\theta$ by maximizing the log-likelihood of the observed data. When $y$ is unobserved, we treat it as a latent variable and marginalize it out. Specifically, we optimize the following expression:
$\underset{\theta}{\text{argmax}}P(\mathbf{S},Y;\theta)\!\!\!\!$ $=$ $\!\!\!\!\underset{\theta}{\text{argmax}}\left[\ \prod_{i=1}^{n_{\ell}}P(\mathbf{s}_{i},y_{i};\theta)\prod_{j=1}^{n_{u}}P(\mathbf{s}_{j};\theta)^{\lambda_{U}}\ \right]$ (1)
$=$ $\!\!\!\!\underset{\theta}{\text{argmax}}\Bigg[\ \underbrace{\sum_{i=1}^{n_{\ell}}\log\left[P_{\theta}(\mathbf{s}_{i}|y_{i})P_{\theta}(y_{i})\right]}_{\shortstack{Labeled Data Log-Likelihood}}+\lambda_{U}\underbrace{\sum_{j=1}^{n_{u}}\log\sum_{k=1}^{K}\left[P_{\theta}(\mathbf{s}_{j}|y_{j}=k)P_{\theta}(y_{j}=k)\right]}_{\shortstack{Unlabeled Data Log-Likelihood}}\ \Bigg]$ (2)
where $\lambda_{U}$ controls the relative weight of the unlabeled data in the likelihood; we fix $\lambda_{U}=1$.
We optimize the above expression using expectation-maximization (EM) (dempster1977maximum, ; zhao2023learning, ), alternating between the E-step, which estimates which mixture component (i.e. ground truth class) each data point belongs to, and the M-step, which estimates the parameters of each component based on the soft component assignments. Algorithm 1 details the EM procedure we use to fit our mixture model.
A technical challenge when fitting densities on bounded domains, as is the case with classifier outputs, is boundary bias, where estimates of density are biased near edges of the domain (jones1993simple, ). To overcome this challenge, we transform probabilistic predictions over $K$ classes to points (“scores") in $\mathbb{R}^{K-1}$ using the additive log-ratio transform, which produces a one-to-one mapping between the two spaces (aitchison_statistical_1982, ; pawlowsky2011compositional, ). Additive log-ratio transforms are a reparameterization trick: they transform the classifier scores from a hard-to-model space — the probability simplex — to an easier-to-model space — unbounded reals. Appendix E.1 provides the expression for the additive log-ratio transform.
Step 2: Estimating classifier performance given $P(y,\mathbf{s})$.
Once we have estimated the parameters $\theta$ for the mixture model, we can use the fitted density to estimate metrics of interest, using the procedure described in Figure 1. By estimating the conditional distribution $P_{\theta}(y|\mathbf{s})$, SSME can measure performance with respect to the expected label of unlabeled examples (where we observe only $\mathbf{s}^{i}$). Specifically, we sample ground truth labels $y$ from the distribution $P_{\theta}(y|\mathbf{s})$ for all unlabeled examples. Labeled examples retain their ground truth labels. We then compute metrics with respect to both the sampled and ground truth labels, and average the estimated metric across samples. SSME can estimate any metric that is a function of classifier scores and ground truth labels; see Algorithm 2 for corresponding pseudocode.
The proposed approach enjoys several properties. By using a semi-supervised mixture model, SSME takes advantage of unlabeled data to estimate $P(y,\mathbf{s})$. SSME also naturally accommodates and benefits from additional classifiers, because they can provide additional information about the ground truth label of an unlabeled example (additional classifiers simply translate to additional dimensions in $\mathbf{s}$). Furthermore, the use of the additive log-ratio transform allows SSME to learn directly from continuous class probabilities for each input, as opposed to discretized labels.
Input: Labeled set $\mathcal{D}_{\ell}=\{(\mathbf{s}^{i},y^{i})\}_{i=1}^{n_{\ell}}$; Unlabeled set $\mathcal{D}_{u}=\{\mathbf{s}^{i}\}_{i=1}^{n_{u}}$; $\mathbf{s}^{i}$ are ALR-transformed classifier scores; Labeled-data weight $\lambda_{U}$; number of classes $K$; total epochs $T$ = 1000
Init: Initialize model parameters $\theta^{0}$;
Initialize class priors $p^{0}(y=k)=\hat{p}(y=k)$ ;
// empirical class distribution
1
2 for $t\leftarrow 0$ to T1T-1 do
3 E-step: compute responsibilities $\gamma_{ik}$;
4  foreach $\mathbf{s}^{(i)}\in\mathcal{D}_{u}$ do
5    for $k\leftarrow 1$ to $K$ do
6       $\gamma_{ik}\leftarrow\frac{p^{t}(y=k)\,P_{\theta^{t}}(\mathbf{s}^{(i)}\mid y=k)}{\sum_{\ell=1}^{K}p^{t}(y=\ell)\,P_{\theta^{t}}(\mathbf{s}^{i}\mid y=\ell)}$;
7
8 foreach $(\mathbf{s}^{i},y^{i})\in\mathcal{D}_{\ell}$ do
    $\gamma_{ik}\leftarrow\mathbf{1}[k=y^{i}]$ ;
     // fix to true label
9
10 M-step: update priors and $\theta$;
11  for $k\leftarrow 1$ to $K$ do
    $N_{k}\leftarrow\sum_{\mathcal{D}_{u}}\gamma_{ik}+\lambda_{U}\sum_{\mathcal{D}_{\ell}}\gamma_{ik}$ ;
     // weighted effective count for class $k$
12     $p^{t+1}(y=k)\leftarrow N_{k}/(n_{u}+\lambda_{U}n_{\ell})$;
13 $\theta^{t+1}\leftarrow\arg\max_{\theta}\sum_{i=1}^{n_{u}+n_{\ell}}w_{i}\sum_{k=1}^{K}\gamma_{ik}\log P_{\theta}(\mathbf{s}^{i}\mid y=k)$,
14  where $w_{i}=\lambda_{U}$ for unlabeled points and $w_{i}=1$ otherwise;
Output: Final parameters $\theta^{T}$ and priors $p^{T}(y)$
Algorithm 1 Estimating $P(y,\mathbf{s})$ using EM
3.1 Implementation
SSME can accommodate multiple ways to parametrize the class-conditional distribution of scores $P(\mathbf{s}\mid y)$ as long as they can be learned in the semi-supervised setting described above. We denote the parameterized distribution as $P_{\theta}(\mathbf{s}\mid y)$. In our experiments, we use a kernel density estimator (KDE) to parameterize $P_{\theta}(\mathbf{s}\mid y)$ and explore alternative parametrizations of $P(\mathbf{s}^{(i)}|y^{(i)})$ in Appendix E.3. Kernel density estimators do not make parametric assumptions about the distributional form of each component: this is useful for modeling distributions of predictions, which can vary widely across tasks and classifiers. The KDE for the $k$th class is parametrized by bandwidth $h$ and the kernel type $\mathcal{K}$. We weight each point by the probability it belongs to the component (i.e., a soft label). When $\mathbf{s}^{i}$ is labeled, $P(y^{i}=k)=1$ for the true label $k$, and when $\mathbf{s}^{i}$ is unlabeled, we compute $P(y^{i}=k|\mathbf{s})$ using the estimated density. For all experiments, we fit use a Gaussian kernel and estimate $h$ using the improved Sheather-Jones algorithm (botev2010kernel, ). We optimize the parameters using EM over 1000 epochs. We initialize component assignments by drawing a label for a given example according to the mean classifier score across the set of classifiers.
3.2 Theoretical analysis
We motivate our approach by deriving bounds for how well a semi-supervised estimation procedure can estimate classifier performance in a setting where classifier scores $\mathbf{s}$ are drawn from a mixture of Gaussians.
Model of $y$ and $\mathbf{s}$.
We consider a stylized binary classification setting drawn from previous work (NEURIPS2023_458fa8ee, ) where classifier scores $\mathbf{s}$ are drawn from a balanced mixture of two Gaussians. For data in the positive class, classifier scores are drawn from a multivariate Gaussian $\mathcal{N}(c,I)$. For data in the negative class, classifier scores are drawn from a multivariate Gaussian $\mathcal{N}(0,I)$. Let $n_{u}$ refer to the number of unlabeled examples, $n_{\ell}$ the number of labeled examples, and $d$ the number of classifiers. Additional classifiers increase the dimensionality of the multivariate Gaussian.
Performance estimation procedure.
Having specified how classifier scores are generated, we analyze how accurately a semi-supervised performance estimation procedure similar to our own can estimate classifier performance. We consider a class of semi-supervised learning algorithms frequently analyzed in prior work, UL+ (NEURIPS2023_458fa8ee, ). UL+ algorithms first use the unlabeled data to identify decision boundaries and then use the labeled data to assign regions to classes. $AUC_{k}$ and $ACC_{k}$ represent ground truth performance of classifier $k$; $\widehat{AUC}_{k}$ and $\widehat{ACC}_{k}$ are performance estimators that use UL+ to estimate a Gaussian mixture model, and use that model to estimate performance. We apply bounds developed by NEURIPS2023_458fa8ee for errors in mixture model estimation to bound errors in performance estimation.
Theorem 1
@@ 3.2 Theoretical analysis
$|AUC_{k}-\widehat{AUC}_{k}|\leq$ $\quad\Phi\big(\frac{\mathbf{c}_{k}}{\sqrt{2}}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{\sqrt{2}}\big)$ (3)
$|ACC_{k}-\widehat{ACC}_{k}|\leq$ $\quad\Phi\big(\frac{\mathbf{c}_{k}}{2}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{2}\big)$ (4)
@@
The expression for $\epsilon_{\mathbf{c}}$ captures its asymptotic dependence on $_{u}, nn_{\ell}$, $d$, and $c$. Smaller $\epsilon_{\mathbf{c}}$ correspond to smaller errors in performance estimation. Section C contains the full proof, which consists of two steps: we first bound the error in estimating the separation of the two Gaussian components $\mathbf{c}||$, and then use the relationship between component separation and classifier performance to bound error in estimates of AUC and accuracy.
Analysis.
Our theoretical bounds have several implications. First, unlabeled data helps: increases in $n_{u}$ reduce both the first term and the second term. This indicates why SSME outperforms baselines that only make use of labeled data: additional unlabeled data improves our ability to accurately estimate components of a mixture model ($\epsilon_{\mathbf{c}}\eqsim O(1/\sqrt{n_{u}})$). Second, bounds tighten as separation between components $\mathbf{c}||$ grows, i.e. when classifiers in the set are more accurate. Third, bounds tighten as the number of classifiers increases, so long as the increase in separation $c||$ outweighs the cost of an additional dimension in $d$; we precisely characterize this setting in Section C.1. This helps explain why SSME outperforms work that makes use of a single classifier, instead of several.
Our theoretical results are derived for a stylized setting frequently studied in prior work ratsaby1995learning ; rigollet2007generalization ; NEURIPS2023_458fa8ee ; frei2022self . While this setting differs from our empirical setting – for example, it uses a simplified semi-supervised learning algorithm – we verify that trends our theory implies hold robustly in both synthetic and real experiments (Sec. 5, D.1), including settings where these assumptions do not hold.